Calculate the amperage of a device with power 460w and voltage 230v


1
Expert's answer
2021-04-23T11:06:44-0400

Explanations & Calculations


  • Those ratings of the device imply that it consumes a power of 460W exactly at an applied voltage of 230V.
  • Lower the applied voltage, lower the power it consumes.


  • With the use of the simple equation "\small P =Vi", the amperage of the device can be calculated.
  • Knowns — "\small V=230V" and "\small \text{Power}\,=P=460W"
  • Then,

"\qquad\qquad
\begin{aligned}
\small i&=\small \frac{P}{V}=\frac{460W}{230V}=\bold{2A} 
\end{aligned}"

  • The device consumes that much of an amperage.